### v1

# WisPaper: Your AI Scholar Search Engine

Abstract
Researchers struggle to efficiently locate and manage relevant literature within the exponentially growing body of scientific publications. We present WisPaper, an intelligent academic retrieval and literature management platform that addresses this challenge through three integrated capabilities: (1) Scholar Search, featuring both quick keyword-based and deep agentic search modes for efficient paper discovery; (2) Library, a customizable knowledge base for systematic literature organization; and (3) AI Feeds, an intelligent recommendation system that automatically delivers relevant new publications based on user interests. Unlike existing academic tools, WisPaper provides a closed-loop workflow that seamlessly connects literature discovery, management, and continuous tracking of research frontiers. Our multilingual and multidisciplinary system significantly reduces the time researchers from diverse backgrounds spend on paper screening and management, enabling them to focus on their core research activities. The platform is publicly accessible and serves researchers across academia and industry.
[…]
[Website]https://wispaper.ai/ footnotetext: Equal Contribution.
Corresponding authors.
1 Introduction
The exponential growth of scientific publications presents unprecedented challenges for researchers attempting to stay current in their fields [5]. Efficiently identifying relevant work and systematically managing findings over time have become increasingly time-consuming and cognitively demanding[12].
Current academic search systems exhibit significant limitations in supporting comprehensive research workflows [8]. Traditional platforms like Google Scholar [^1] https://scholar.google.com and Semantic Scholar [^2] https://www.semanticscholar.org/ rely primarily on keyword-based retrieval, which struggles with conceptual queries and semantic understanding. Recent advances have leveraged large language models (LLMs) [15, 3, 20] for query reformulation and expansion [10, 2, 14, 16], improving initial query understanding. Yet accurate literature discovery fundamentally requires scholarly judgment about what papers actually contribute, a capability that current retrieve-then-rank pipelines cannot provide as they lack mechanisms for deep reasoning over paper content and research context.
Beyond search, existing tools address only isolated aspects of the research process. Citation network visualizations such as Connected Papers [^3] https://www.connectedpapers.com/ reveal paper relationships but lack integrated search and management functions. Reference managers like Zotero [^4] https://www.zotero.org/ and Mendeley [^5] https://www.mendeley.com/ excel at organization yet provide minimal discovery capabilities and require manual curation. Critically, no existing solution bridges the entire research lifecycle from intent-driven discovery through systematic management to continuous frontier tracking, forcing researchers to stitch together fragmented tools with manual effort at each transition point.
We introduce WisPaper, a comprehensive platform that addresses these challenges through three interconnected modules: (1) Scholar Search provides dual-mode retrieval: Quick Search for rapid keyword matching and Deep Search that employs agentic reasoning to evaluate papers’ actual contributions for complex conceptual queries, balancing speed and depth according to user needs. (2) Library enables systematic, custo-
mizable organization of both search results and uploaded documents, supporting long-term knowledge accumulation and continuous updating. (3) AI Feeds automatically recommends highly relevant new papers based on user-defined research interests and library content, enabling researchers to maintain awareness of frontier developments without manual monitoring.
These modules form a closed-loop workflow: researchers search for papers, import and organize them in their Library, while AI Feeds continuously supplements relevant new literature. This integration supports the entire research lifecycle from initial exploration to sustained knowledge management.
Figure 1: WisPaper Knowledge Loop
WisPaper serves diverse user groups including academic researchers, graduate students, and industrial research and development teams across multiple disciplines and languages. By reducing time spent on paper screening and management, the platform enables researchers to allocate more cognitive resources to their core intellectual work. Our contributions are threefold:
(1) A unified platform integrating the entire academic workflow from literature discovery through systematic organization to continuous frontier monitoring, eliminating the need for fragmented tools.
(2) A dual-mode search architecture that combines rapid keyword retrieval with deep semantic reasoning over research context, accommodating diverse search requirements.
(3) A personalized recommendation engine that constructs dynamic user profiles from library curation patterns and explicit interests, reducing manual screening overhead while maintaining research awareness.
2 WisPaper
In this section, we detail the three core modules of WisPaper and their interconnections.
2.1 Scholar Search Module
Scholar Search serves as the primary entry point for literature discovery in WisPaper, offering two complementary search modes that address fundamentally different research scenarios.
2.1.1 Quick Search: Millisecond-Level Responsive Literature Retrieval
Quick Search is is built on efficient keyword retrieval for scenarios where researchers have concrete search targets. When a user has a specific paper title in mind, knows the exact author they want to follow, or needs to quickly scan a well-defined topic, Quick Search provides millisecond-level responses that feel instantaneous.
Each search result presents a complete picture of the paper: beyond the basic title and author information, users can immediately see institutional affiliations, publication dates, venue prestige, citation counts that indicate impact, and abstract snippets that provide context for quick relevance judgments. The interface supports rich interactions—users can access direct citation formatting, express preferences through like and dislike actions, and add papers with available PDFs directly into their Library with one click. More importantly, the system maintains a complete history of all search queries, allowing researchers to re-execute past queries or use them as starting points for refined searches.
2.1.2 Deep Search: Agent-Powered Intelligent Literature Exploration
Deep Search represents a fundamentally different paradigm: an AI agent that conducts intelligent literature exploration on behalf of researchers. Rather than reducing research questions to keywords, the system deploys an autonomous agent (WisModel) that reasons about information needs and orchestrates retrieval tools to discover relevant papers.
Agent-Powered Workflow. The exploration process consists of two key stages orchestrated by WisModel:
  • Query Understanding Stage: The agent interprets the natural language query, performs query decomposition to identify essential aspects, and constructs structured validation criteria (detailed prompt templates are provided in Appendix Table 7).
  • Paper Validation Stage: For each candidate paper, the agent evaluates whether it satisfies the constructed criteria, extracting relevant evidence and producing explicit classifications using a systematic assessment taxonomy (Appendix Table 10). The complete validation prompt structure is shown in Appendix Tables 8 and 9.
Example Workflow. When asked “How do transformer models handle long sequences efficiently?”, WisModel identifies key aspects (transformer architecture, sequence length challenge, efficiency methods), generates targeted search queries, and establishes validation criteria. It then evaluates each retrieved paper against these criteria, classifying results as “Perfect” (meeting all criteria), “Partial” (addressing some aspects), or “No” (failing requirements).
Collaborative Refinement. A distinguishing feature is the agent’s collaborative refinement capability. Researchers can examine and modify the agent-generated criteria to better reflect their information needs, then re-execute the search with refined criteria, progressively narrowing toward their target papers.
2.2 Library Module
Library acts as the central component of WisPaper, evolving the platform from a search tool into a comprehensive knowledge management environment. The design philosophy recognizes that effective research requires structured organization, convenient retrieval, and seamless integration with the reading process.
Organization and Retrieval: Library implements a hierarchical folder system with unlimited nesting depth, allowing researchers to construct taxonomies aligned with their domain structure. Cross-library search addresses the scenario where researchers recall paper content but not storage location, supporting both comprehensive library queries and subtree-constrained searches.
Paper Import and Metadata Extraction: Papers enter the Library through three mechanisms: direct PDF upload, one-click import from Scholar Search results, or integration from AI Feeds recommendations. Upon import, the system automatically extracts metadata including title, authors, affiliations, abstract, publication venue, date, DOI, and references through PDF structure analysis.
Reading Environment and Ecosystem Integration: The integrated PDF reader extends Library beyond storage into active research support. AI-generated structured summaries enable rapid paper assessment, while community annotations surface peer insights. Personal annotation tools provide highlighting and note-taking with full-text search across notes. Export capabilities support standard bibliographic formats (BibTeX, RIS, EndNote) for manuscript preparation, while synchronization plugins for tools like Zotero and Mendeley enable researchers to maintain WisPaper libraries while using familiar citation workflows.
2.3 AI Feeds Module
AI Feeds addresses one of the most cognitively demanding aspects of active research: staying current with rapidly evolving literature. In fast-moving fields, researchers spend 1-2 hours daily manually scanning journals, conferences, and preprint servers.
AI Feeds automates literature monitoring through dual-layer filtering. Users first select broad research areas (e.g., ArXiv categories such as cs.CL for Computation and Language). The system retrieves all papers in these categories since last login, typically 50-100 papers daily per category. Second-layer filtering applies semantic matching against user-defined interests: specific authors, institutions, and topical descriptions. For example, selecting cs.CL with interests in “vision-language models from Google/DeepMind” filters 50-100 daily papers to 3-5 highly relevant matches.
The system learns implicitly from user behavior to refine future filtering. Papers that users add to Library signal strong relevance, while consistently dismissed papers suggest the interest model needs adjustment. This feedback loop operates automatically, gradually improving filtering precision.
2.4 Module Interconnection
WisPaper’s power emerges from the seamless integration of its three modules into a unified research workflow. At the core is a shared paper database that ensures any paper discovered, filtered, or stored in one module is immediately accessible across all others, eliminating redundancy and maintaining consistency.
The modules continuously exchange information to enhance each other’s functionality. Library contents inform AI Feeds’ filtering algorithms, while Scholar Search patterns influence which papers surface in AI Feeds. Papers flow effortlessly between modules with minimal interaction—a single click moves content from discovery to collection while preserving metadata and context.
A typical research session illustrates this integration: check AI Feeds for new publications → add relevant papers to Library → encounter an unfamiliar concept while reading → launch Deep Search from within the reader → organize discovered papers into project folders. Throughout, researchers remain in a single environment without manual data transfer.
3 Deep Search: WisModel Training Pipeline
Technical Implementation Overview. WisPaper’s three modules employ different technical approaches suited to their respective objectives. Quick Search builds upon established keyword-based retrieval systems optimized for millisecond-level response times, while AI Feeds adapts conventional recommendation and filtering frameworks to the academic literature domain. Our primary algorithmic contribution lies in Deep Search’s agent-powered exploration capability, which requires novel training procedures to enable the sophisticated reasoning and validation behaviors described in Section 2.1.2. This section details the training methodology for WisModel, the AI agent that powers Deep Search.
3.1 Training Objectives
WisModel is trained to accomplish two primary objectives that mirror the key stages of agent-powered literature exploration described in Section 2.1.2:
Objective 1: Query Understanding and Criteria Generation.
Given a natural language research query $q$ and current timestamp $t$, the model must generate a set of Boolean search queries $\mathcal{S}=\{s_{1},s_{2},\ldots,s_{k}\}$ (where $k\in[2,4]$) and a set of validation criteria $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{m}\}$ (where $m\in[1,4]$). Each criterion $c_{i}$ consists of a type (e.g., task, methodology, dataset), a descriptive name, a detailed description, and an importance weight $w_{i}$ such that $\Sigma_{i=1}^{m}w_{i}=1$. The complete prompt template is provided in Appendix Table 7.
Objective 2: Paper-Criteria Matching Assessment.
Given a research query $q$, a set of validation criteria $\mathcal{C}$, and paper metadata $p$ (including title, authors, abstract, venue, publication date, and citation count), the model must produce an assessment $a_{i}\in\{\text{support},\text{somewhat\_support},\text{reject},\text{insufficient\_information}\}$ for each criterion $c_{i}\in\mathcal{C}$. The assessment taxonomy is detailed in Appendix Table 10, with prompt templates in Appendix Tables 8 and 9. For each assessment, the model must provide a natural language explanation and extract supporting evidence from the paper metadata.
To provide concrete understanding of the two training objectives, we present two representative examples that demonstrate WisModel’s reasoning process. Table 5 shows a complete query-to-criteria workflow for Objective 1, while Table 6 demonstrates WisModel’s nuanced validation capability for Objective 2—particularly its ability to distinguish between full and partial support, a challenge that proves difficult for baseline models (as quantitatively validated in Section 4.3).
3.2 Training Procedure
Stage 1: Supervised Fine-Tuning.
We initialize WisModel from a pre-trained language model and perform supervised fine-tuning (SFT) on both objectives simultaneously. For Objective 1, the model learns to generate structured JSON outputs containing search queries and criteria. For Objective 2, the model learns to produce detailed assessment objects with criterion-level verdicts, explanations, and evidence extraction. This cold-start phase establishes the basic capability to understand academic queries and reason about paper relevance.
Stage 2: Group Relative Policy Optimization.
Following SFT, we apply Group Relative Policy Optimization (GRPO) [18] specifically to Objective 2 to refine the model’s validation capabilities. Given a prompt $x$ (comprising query, criteria, and paper metadata) with ground truth assessment $a^{*}\in$ {support, somewhat_support, reject, insufficient_information}, the reward function evaluates whether the model’s predicted assessment $a$ aligns with the expert annotation:
$r\text{(}a,a^{*}\text{)}=\mathbb{1}\text{[}a=a^{*}\text{]}=\begin{cases}1&\text{if }a=a^{*}\\ 0&\text{otherwise}\end{cases}$ (1)
GRPO samples $G$ output candidates $\{o_{1},\ldots,o_{G}\}$ from the current policy $\pi_{\theta}$ and computes their respective rewards $\{r_{1},\ldots,r_{G}\}$. We maximize the following objective:
[…]
where $A_{i,t}=\text{(}r_{i}-\text{mean}\text{(}\mathbf{r}\text{))}\text{/std}\text{(}\mathbf{r}\text{)}$ is the advantage for each token.
4 Evaluation
4.1 Evaluation Setup
We conduct comprehensive evaluations on the two critical components underlying Deep Search (Section 2.1.2): (1) query understanding and criteria generation (Objective 1 in Section 3.1), and (2) paper-criteria matching accuracy (Objective 2 in Section 3.1). These evaluations assess WisModel’s ability to interpret user intent and accurately identify relevant academic literature.
Data Collection and Annotation: We recruited domain-expert PhD students to annotate real-world queries spanning 10 academic disciplines. Our dataset comprises 2,777 authentic queries in both Chinese and English, covering diverse research areas including computer science, biology, physics, economics, and social sciences. For each query, annotators provided:
  • Search queries: Optimized keyword combinations for academic database retrieval.
  • Criteria: Specific conditions to determine whether a paper falls within the target scope.
The distribution statistics are shown in Figure 2, and representative query examples with their corresponding criteria are presented in Table 1.
Example Query-Criteria Pairs. Table 1 illustrates the diversity of search intents across disciplines. For instance, a Computer Science query asks “Are there any papers that investigate the relationship between in-context learning and inference-time scaling?”, with criteria requiring papers to investigate both in-context learning methods and inference-time scaling methods. A Medicine query seeks “Find the ELN guideline that recommends ruxolitinib for patients with myelofibrosis”, with criteria specifying both guideline type (ELN) and treatment recommendation content. These examples demonstrate how WisModel must decompose complex queries into verifiable, independent criteria.
Domain Query Example Criteria Example
[…]
Medicine Find the ELN guideline that recommends ruxolitinib for patients with myelofibrosis. 1. The paper is an ELN (European LeukemiaNet) guideline or consensus statement. 2. The paper recommends or includes ruxolitinib as a treatment option for patients with myelofibrosis.
[…]
Psychology Find related papers about adults use physical experiences to understand and reason about spatial concepts. 1. The paper investigates spatial reasoning or understanding in adults. 2. The study uses physical experience or interaction as a method to explore spatial concepts.
[…]
Physics Find papers that describe how to implement three position and three velocity into semi analytical models for galaxy/halo formation. 1. The paper describes or implements a method for incorporating three position and three velocity into semi-analytical models. 2. The study focuses on galaxy or halo formation processes.
[…]
Neuro Science Find papers about longtern Marajuan use and the brain 1. The paper investigates the effects of long-term marijuana use. 2. The study examines brain-related outcomes or changes.
[…]
Chemistry Paper describing the contribution of chemical oscillations to the understanding of the dynamism of life phenomena. 1. The paper discusses the role or contribution of chemical oscillations to the understanding of life phenomena. 2. The paper addresses the dynamism or dynamic behavior of life phenomena.
Table 1: Real-world queries and annotated criteria across 10 academic disciplines. Examples demonstrate the variety of search intents and the multiple criteria users specify when seeking relevant academic papers in different domains.
Figure 2: Distribution statistics of the evaluation dataset across 10 academic disciplines. The left panel shows the average query length in tokens for each discipline. The right panel displays the average number of criteria per query. These statistics indicate the complexity of search requirements across different fields.
4.2 Query Understanding and Criteria Generation
Task Description. This experiment evaluates whether WisModel’s generated outputs (search queries and criteria following the format in Appendix Table 7) align with human expert annotations. This corresponds to the first stage of Deep Search’s agent workflow (Section 2.1.2).
Evaluation Metrics: We employ multiple complementary metrics:
  • Semantic Similarity: Cosine similarity using OpenAI’s text-embedding-3-large embeddings [^6] https://platform.openai.com/docs/guides/embeddings, measuring semantic alignment between generated and reference texts.
  • ROUGE scores: ROUGE-1, ROUGE-2, and ROUGE-L for n-gram overlap evaluation.
  • BLEU: Precision-focused metric for exact phrase matching.
  • Length Ratio: Percentage comparing generated text length to reference length (optimal: 100%)
Baseline Models: We compare WisModel against state-of-the-art language models:
  • GPT-5.1: OpenAI’s latest flagship model with adaptive reasoning mechanisms.
  • GPT-4o: OpenAI’s widely-adopted commercial model with multimodal capabilities.
  • Qwen-Max: Alibaba’s high-performance model with trillion-scale MoE architecture.
  • GLM-4-Flash and GLM-4.6: Zhipu AI’s models, with GLM-4.6 (355B parameters) leading in coding tasks.
  • DeepSeek-V3.2-Exp: Leading open-source alternative with strong reasoning performance.
Results: Table 2 presents the comprehensive evaluation results. Key findings include:
(1) WisModel establishes new state-of-the-art: WisModel achieves dominant performance across all metrics, with Semantic Similarity of 94.8%, ROUGE-L of 67.7%, and BLEU of 39.8%. This represents substantial improvements of 4.8%, 15.1%, and 18.3% respectively over the second-best model.
(2) DeepSeek-V3.2-Exp emerges as the breakthrough open-source model: DeepSeek-V3.2-Exp demonstrates remarkable performance, achieving 90.2% semantic similarity—surpassing several commercial models. This marks a significant milestone where open-source models match or exceed proprietary alternatives in specialized academic tasks.
(3) GPT-4o maintains reliability: GPT-4o delivers consistent performance (91.3% semantic similarity, 52.6% ROUGE-L, 21.5% BLEU), making it a dependable choice for production deployments.
(4) Model-specific trade-offs emerge: While GPT-5.1 achieves strong semantic understanding (87.0%), it struggles with exact phrase reproduction (BLEU 13.2%). This suggests that newer models do not universally improve performance across all evaluation dimensions.
Model Semantic Similarity ROUGE-1 ROUGE-2 ROUGE-L BLEU Length Ratio
Qwen-Max 78.1 43.2 23.1 35.8 11.8 168.9
GPT-4o 91.3 64.0 39.4 52.6 21.5 142.2
GPT-5 87.0 53.8 27.6 41.8 13.2 163.3
GLM-4-Flash 82.2 50.0 25.8 42.1 9.9 167.1
GLM-4.6 84.8 55.5 30.2 44.5 14.4 168.1
DeepSeek-V3.2-Exp 90.2 59.3 32.4 48.0 14.4 153.5
WisModel 94.8 74.9 54.4 67.7 39.8 98.2
Table 2: Model performance on query understanding and criteria generation task. Evaluation metrics include semantic similarity, ROUGE scores, BLEU, and length ratio. WisModel outperforms all baseline models across all dimensions.
Application Recommendations: Based on our findings:
  • Maximum Quality: WisModel achieves unmatched performance, suitable for mission-critical applications.
  • Cost-Efficiency: DeepSeek-V3.2-Exp offers exceptional value as an open-source alternative.
  • Exact Matching: For tasks requiring precise phrase reproduction, WisModel (BLEU 39.8%) significantly outperforms alternatives.
  • Enterprise Deployments: GPT-4o provides balanced performance with comprehensive ecosystem support.
4.3 Paper-Criteria Matching
Task Description: This experiment evaluates WisModel’s ability to accurately determine whether papers satisfy specified criteria, corresponding to the second stage of Deep Search (Section 2.1.2). Given the 2,777 queries and their 5,879 human-annotated criteria from Section 5.2, we evaluate matching accuracy across the four assessment categories defined in our taxonomy (Appendix Table 10). Domain-expert PhD students labeled each paper-criteria pair using these categories. The complete prompt structure for this task is provided in Appendix Tables 8 and 9.
Results: Table 3 presents the matching accuracy for each category. Key findings include:
(1) WisModel demonstrates overwhelming superiority: WisModel achieves 93.70% overall accuracy compared to the next best model at 73.23%, with strong performance across all four categories (90.64% insufficient information, 94.54% reject, 91.82% somewhat support, 94.38% support).
(2) Category-specific baseline performance varies dramatically:
  • Support Recognition: Gemini-3-Pro excels at identifying fully supportive papers (91.1%).
  • Reject Detection: Qwen3-Max achieves the highest baseline rejection accuracy (72.0%).
  • Insufficient Information: Both Gemini-3-Pro and GPT-5.1 exceed 64% accuracy.
  • Somewhat Support - A Shared Weakness: Baseline models struggle significantly, with Gemini-3-Pro achieving only 15.9%—a dramatic drop from its 91.1% support accuracy.
(3) The “partial support” challenge: The poor performance on “somewhat support” reveals a critical limitation of existing models—they lack the nuanced judgment required for gray-area cases where papers partially address criteria. This represents a key capability that distinguishes WisModel from current alternatives.
Model Criteria Categories Overall Accuracy
insufficient information reject somewhat support support
GPT-5.1 64.30 63.10 31.40 85.40 70.81
Claude-Sonnet-4.5 46.00 66.50 33.30 87.00 70.62
Qwen3-Max 40.80 72.00 44.20 87.20 72.82
DeepSeek-V3.2 57.90 49.20 45.00 87.00 66.82
Gemini3-Pro 67.40 66.80 15.90 91.10 73.23
WisModel 90.64 94.54 91.82 94.38 93.70
Table 3: Model accuracy on paper-criteria matching task. Performance is measured across four classification categories and overall accuracy. WisModel significantly outperforms all baseline models, achieving 93.70% overall accuracy compared to the next best model at 73.23%.
Application Recommendations for Baselines: While WisModel dominates in academic paper retrieval scenarios, users without access may consider task-specific alternatives shown in Table 4. For identifying “full support” cases, Gemini-3-Pro achieves the highest match rate (91%). For rejection detection, Qwen3-Max leads with 72% accuracy. For balanced overall performance, Qwen3-Max shows no extreme weaknesses across categories.
Application Scenario Recommended Model Rationale
Identifying “full support” Gemini-3-Pro Highest support match rate (91%)
Identifying “rejection” Qwen3-Max Highest reject match rate (72%)
Assessing information sufficiency Gemini-3-Pro / GPT-5.1 Both exceed 64% accuracy
Detecting partial support DeepSeek-V3.2 / Qwen3-Max $\sim$45% accuracy, though suboptimal
Balanced overall performance Qwen3-Max No extreme weaknesses across categories
Table 4: Model Recommendation for Different Application Scenarios
5 Related Work
5.1 Academic Retrieval and Discovery
Academic search is a cornerstone of scientific research, with mainstream keyword-based systems such as Google Scholar [^7] https://scholar.google.com/ and PubMed [^8] https://pubmed.ncbi.nlm.nih.gov/ supporting global researchers. However, these tools often fall short in handling complex conceptual queries that demand semantic comprehension, as they rely primarily on surface-level text matching rather than deep insight into research intent [8]. To address this limitation, recent studies have explored query reformulation and expansion techniques, yet these approaches still fail to break through the constraints of literal matching [10, 2, 14, 16]. Subsequent advancements have introduced more sophisticated features: Semantic Scholar [^9] https://www.semanticscholar.org/ integrated citation-based metrics and paper recommendation functions, while Connected Papers [^10] https://www.connectedpapers.com/ enabled visual exploration of citation networks. These tools focus on relationship discovery but lack support for comprehensive literature management.
More recently, LLM-powered agents have brought new breakthroughs to academic search. LitSearch [1] and ResearchArena [11] have begun benchmarking LLM capabilities in this domain, verifying their potential to enhance retrieval performance. PaSa developed an autonomous agent that invokes search tools, parses papers, and navigates citation networks to address complex scholarly queries[9]. PaSa outperforms baselines like Google Scholar and ChatGPT in recall metrics. Nevertheless, PaSa remains confined to the search phase and does not address post-retrieval needs such as literature organization and frontier tracking.
5.2 LLM Agents for Research Assistance
Recent efforts have shifted toward integrating discrete research tasks into cohesive workflows. Specialized systems like ChemCrow [6] and Coscientist [7] automate chemistry experimentation, while virtual laboratory frameworks [19] assist in wet-lab design. However, these tools are domain-specific and do not support cross-disciplinary literature management. End-to-end automation systems such as AI Scientist [13] and ResearchAgent [4] aim to cover ideation, experimentation, and paper writing. While innovative, their fully autonomous design limits user control and positions them as replacements rather than assistants for human researchers. Agent Laboratory [17] expands the automated workflow to include literature review and report writing but prioritizes generating new research outputs over managing existing literature.
6 Conclusion
In this paper, we introduced WisPaper, an integrated platform that unifies intelligent literature discovery, systematic knowledge management, and personalized frontier tracking. By combining fast keyword retrieval, agent-based deep semantic search, and a structured Library with AI-driven updates, WisPaper provides a closed-loop workflow that supports the full research cycle from exploration to long-term curation. Our large-scale evaluation demonstrates that WisModel achieves state-of-the-art performance in both query interpretation and paper–criteria matching, substantially outperforming strong commercial and open-source baselines. We believe WisPaper offers a practical step toward more efficient, reasoning-centered academic workflows
Acknowledgments
We sincerely thank the following individuals for their valuable contributions to the data annotation process: Jiaqi Du (Nanjing University), Zixuan Wang (Fudan University), Xuetong Zhue (Tsinghua University), Tianjiao Xue (Tsinghua University), Shuyue Zhang (Sydney University), Feng Cui (Fudan University), Yonghao Liu (Zhejiang University), Yehao Zhang (Fudan University). Their dedicated efforts and high-quality work were essential to this research.
Appendix
Component Content
User Query I am looking for articles that show how embedding psychologists into subspecialty medical teams reduce medical provider’s level of burnout.”
Generated Query 1: embedded psychologist” AND subspecialty medical team” AND burnout
Search Queries Query 2: psychologist integration” AND “medical team” burnout
Query 3: psychologist embedded medical team burnout
Generated Criteria Criterion 1 (weight: 0.4, type: task)
   Name: Embedded psychologists
   Description: The paper examines the integration or embedding of psychologists.
Criterion 2 (weight: 0.3, type: task)
   Name: Subspecialty medical teams
   Description: The paper is within subspecialty medical teams.
Criterion 3 (weight: 0.3, type: task)
   Name: Impact on burnout reduction
   Description: The paper presents evidence or findings on the effect of embedded psychologists on reducing burnout among medical providers.
Table 5: End-to-end example of WisModel’s query understanding and criteria generation (Objective 1). Given a natural language query about embedding psychologists in medical teams, WisModel performs query decomposition, generates three targeted Boolean search queries, and constructs three weighted validation criteria that operationalize the user’s information need into verifiable checkpoints.
Component Content
User Query Find a paper that argues remote workers who received limited feedback from supervisors experienced lower levels of motivation and job satisfaction and suggests that regular, constructive feedback is critical for maintaining engagement and enhancing employees’ psychological well-being in remote work environments.”
Validation Criteria Criterion 1: The paper examines the effects of limited feedback from supervisors on remote workers’ motivation and job satisfaction.
Criterion 2: The paper argues that regular, constructive feedback is critical for maintaining engagement and enhancing employees’ psychological well-being in remote work environments.
Paper Metadata Title: Organization of remote work in the context of digitalization
Authors: Anna Sheveleva, Evgeny Rogov
Venue: Developmental Psychology (Journal)
Year: 2021
Abstract (excerpt): …Individual-personal effects are connected with professional’s personality transformation under influence environment digitalization: a change emotional side experiencing lack information about colleagues’ work, feedback their role overall result…Organizational managerial associated implementation regulation, control assessment, staff motivation, forms employment, membership commitment, job satisfaction…”
WisModel Criterion 1: Support
Assessment    Explanation: The abstract discusses job satisfaction and motivation in the context of remote work, directly addressing the criterion.
   Evidence: Organizational managerial associated…staff motivation…job satisfaction, trust teams…”
Criterion 2: Somewhat Support
   Explanation: The abstract mentions a lack of feedback (albeit not explicitly from supervisors), but does not explicitly argue that constructive feedback is critical for maintaining engagement or enhancing psychological well-being.
   Evidence: “…experiencing lack information about colleagues’ work, feedback their role overall result…”
Summary The paper analyzes remote work organization in a digitalized context, exploring individual, social, and managerial effects such as lack of feedback and its potential impact on job satisfaction. However, it does not clearly argue that regular, constructive feedback is essential to maintain engagement or improve psychological well-being in remote work environments.
Table 6: Example of WisModel’s paper-criteria validation (Objective 2) demonstrating nuanced assessment capabilities. Given a paper on remote work organization, WisModel correctly identifies support for Criterion 1 (clear evidence present in abstract) and somewhat_support for Criterion 2 (indirect evidence only). This ability to distinguish between full and partial support represents a key strength—baseline models achieve only 15.9–45.0% accuracy on somewhat_support” cases (Table 3), while WisModel reaches 91.82%.
System Prompt:
Your name is WisModel, an expert in academic search and literature screening. Your job is to:
1) Generate 2-4 Google Scholar search queries ("search_queries").
2) Generate 1-4 executable, standalone screening criteria ("criteria"), each an independent rule.
User Prompt:
Current time: {timestamp}.
User query: {user_query}
Expected Output Format (JSON):
"search_queries": [
"<Boolean search expression 1>",
"<Boolean search expression 2>",
"criteria": [
"type": "<task|method|dataset|metric|etc.>",
"name": "<Short criterion name>",
"description": "<Detailed criterion description>",
"weight": <float between 0 and 1>
Table 7: Prompt template for query decomposition and criteria generation (Objective 1).
System Prompt:
You are WisModel, a meticulous academic content auditor. Your task is to act as an academic expert and strictly follow a set of rules to verify if a given academic paper (paper_info) aligns with a set of criteria derived from a user’s query.
User Prompt:
Current time: {timestamp}
Original user query: {user_query}
**Validation criteria:**
<criteria>
<criterion_1>{criterion_1_description}</criterion_1>
<criterion_2>{criterion_2_description}</criterion_2>
</criteria>
**Paper details for validation:**
<paper_info>
<title>{paper_title}</title>
<authors>{paper_authors}</authors>
<affiliations>{paper_affiliations}</affiliations>
<conference_journal>{venue_name}</conference_journal>
<conference_journal_type>{venue_type}</conference_journal_type>
<research_field>{research_fields}</research_field>
<doi>{paper_doi}</doi>
<publication_date>{publication_date}</publication_date>
<abstract>{paper_abstract}</abstract>
<citation_count>{citation_count}</citation_count>
<source_url>{paper_url}</source_url>
</paper_info>
**Your Task:**
Based on the rules provided in your instructions, you must perform a rigorous, step-by-step validation and generate a single JSON object as your response. Write all text fields (explanation, summary) in **English**.
Table 8: Prompt template for paper-criteria alignment assessment (Objective 2).
Expected Output Format (JSON):
"criteria_assessment": [
"criterion_id": "<criterion_id>",
"assessment": "<support|somewhat_support|reject|insufficient_information>",
"explanation": "<Natural language explanation for the assessment>",
"evidence": [
"source": "<title|abstract|authors|etc.>",
"text": "<Extracted text supporting the assessment>"
"summary": "<Overall assessment summary explaining whether and how well the paper addresses the original query>"
Table 9: Expected output format for paper validation (Objective 2).
Category Definition and Examples
Support The paper contains clear, direct evidence that confirms or fully satisfies the criterion.
Example: Criterion requires "model parameters exceeding 1 billion"; paper states "our model contains 1.2 billion parameters."
Reject The paper contradicts the criterion or is fundamentally irrelevant. Two subcategories:
(a) Explicit Contradiction: Direct evidence negating the criterion (e.g., criterion requires "asynchronous training"; paper uses only "synchronous training").
(b) Foundational Irrelevance: Domain/topic mismatch makes criterion inapplicable (e.g., criterion about "Chinese economic policy" applied to "Roman Empire agriculture").
Somewhat Support The paper partially relates to the criterion but provides incomplete, indirect, or limited evidence.
Example: Criterion requires testing on "ImageNet and COCO datasets"; paper only reports ImageNet results.
Insufficient Information Available metadata (title, abstract, publication date) provides neither supporting nor contradicting evidence; relationship cannot be determined.
Example: Criterion asks about "experiments on physical robots"; abstract mentions neither platform details nor experimental setup.
Table 10: Assessment taxonomy and decision rules for paper validation.